Assume that you hold a well-diversified portfolio that has an expected return of 11.0% and a beta of 1.20. You are in the process of buying 1,000 shares of Alpha Corp at $15 a share and adding it to your portfolio. Alpha has an expected return of 21.5% and a beta of 1.70. The total value of your current portfolio is $85,000. What will the expected return and beta on the portfolio be after the purchase of the Alpha stock? R ound your answer to two decimal places.

a.

12.41%; 1.56

b.

13.98%; 1.26

c.

12.05%; 1.25

d.

12.58%; 1.28

e.

9.40%; 1.34
Answer #1

The percentage of current portfolio in the new portfolio will be = 85000/(85000 + 15000) = 85% and that of alpha corp will be 15%. Hence, beta and expected returns are the weighted average of the portfolio and the shares.

Beta = 1.2 x 0.85 + 1.7 x 0.15 = 1.275

Expected Return = 11 x 0.85 + 21.5 x 0.15 = 12.575%.

Option D is correct.